A statistics student reported getting 9207 heads and 8743 tails out of 17950 coin tosses. is this evidence that the coin is fair? coclysion at the 1percent level
Answer #1

for a coin to be fair ; proportion of head should be 0.5.

null Hypothesis:              Ho:   p= 0.500
alternate Hypothesis:    Ha: p ≠ 0.500
for 0.01 level and two tailed test , critical z= 2.576 (from excel:normsinv(0.005)
Decision rule :   reject Ho if absolute value of test statistic |z|>2.576
sample success x   = 9207
sample size          n    = 17950
std error σp =√(p*(1-p)/n) = 0.0037
sample prop p̂ = x/n=9207/17950= 0.5129
z =(p̂-p)/σp=(0.513-0.5)/0.004= 3.46
since test statistic falls in rejection region we reject null hypothesis
we have sufficient evidence at 0.01 level of significance to conclude that coin is not fair